Can someone please help! I had these multiple choice questions for homework and get them wrong.
(5 quare root of 6-4)(3 square root of 6 - 1)

Answer by checkley75(3666)

THE ONLY SOLUTION YOU OFFER THAT IS CLOSE TO MINE IS
B) 86-17SQRT6=86-17*2.45=86-41.64=44.36
MAYBE I HAVE MISREPRESENTED THE PROBLEM!!!!
-------------------------------
(5SQRT[6-4])(3SQRT[6-1])
5(SQRT2)3(SQRT5)
5*3(SQRT2)(SQRT5)
15*1.414*2.236
47.426